Project

General

Profile

NDOS CommissionMtng 20121001

Attending: Xuebing, Zukai, Martin, Karen, Athans, Peter, Andrew, Kurt, Mat

Weekly Topics:
-Update on FarDet snout test stand
-New Control Room layout
-Zukai NDOS noise study
-Control room computer updates

UPDATES:
-Detector Update
-DataCheck Update
-Installation Plans for the week
-DAQ Update and Plans for the week
-DCS Update and Plans for the week
-AOB

Test Stand-
Karen- Karen informed AD we're pulling back techs for NOvA
- Will work with Erik Vorin to assem
- Work will
- 2 sets of 16 with PVC to make a fake block
- hanging shelf near the MC
- 16 horizontals and 16 verticals.
- Argonne get 1 ft chunks. try to use these.
- Karen cannot ask tech - what's the effort code. (5 weeks of 2 techs)
-Erik Voirin (need an estimate on the utilities)
-Karen estimates (2 FTE weeks for the structure)
-Karol volunteered to send a student

Leon- APDs available ~ Nov. 11
Peter - Detector ID situation is resolved

Control Room update procedure:
- How do we handle these updates.
- Nightly builds
- Rolling updates of daq machines is not expressed
- I'll circulate this document.
Peter - ask if there is another group, andrew no
I'll talk to Bonnie.

Zukai- doc 7913
-HV setting (lets set that up) (Use proper HV for his upcoming noise study)

Jan - No db update

DAQ Software- Jon is looking in RC issues
DB prod access is denied; DCS is logging locally current

Dev DB migrated wednesday

Power supply monitoring - is writing to local disk

Athans - Power supply GUI writing locally for 20. Jon will work with them to write to DB.

DDT - Want to write data to buffer node. (Hold off til after Jon Paley finishes)